OpenStudy (anonymous):

Help I've been stuck on this forever! Divide (9x^4y^3 + 3x^3y^2 - 6x^2y -12x^2y^4) by -3x2y.

OpenStudy (anonymous):

1. Simplifying -5(6x + -7y + 1) = 0

OpenStudy (anonymous):

Solution for -5(6x-7y+1)= equation: Simplifying -5(6x + -7y + 1) = 0 Reorder the terms: -5(1 + 6x + -7y) = 0 (1 * -5 + 6x * -5 + -7y * -5) = 0 (-5 + -30x + 35y) = 0 Solving -5 + -30x + 35y = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '5' to each side of the equation. -5 + -30x + 5 + 35y = 0 + 5 Reorder the terms: -5 + 5 + -30x + 35y = 0 + 5 Combine like terms: -5 + 5 = 0 0 + -30x + 35y = 0 + 5 -30x + 35y = 0 + 5 Combine like terms: 0 + 5 = 5 -30x + 35y = 5 Add '-35y' to each side of the equation. -30x + 35y + -35y = 5 + -35y Combine like terms: 35y + -35y = 0 -30x + 0 = 5 + -35y -30x = 5 + -35y Divide each side by '-30'. x = -0.1666666667 + 1.166666667y Simplifying x = -0.1666666667 + 1.166666667y

OpenStudy (anonymous):

I am confused, that's an example right?

OpenStudy (wmj259):

He actually solved the problem in the most simplest form, I just skimmed over it. long post :).

OpenStudy (anonymous):

the answer is -0.000111111111133

OpenStudy (wmj259):

\[\frac{ 9x^4y^3+3x^3y^2-6x^2y-12x^2y^4}{ -3x2y }\]

OpenStudy (anonymous):

Okay So first I divide -3x2y and 9x4y3 ?

OpenStudy (anonymous):

hope that i helped

OpenStudy (wmj259):

@randomrefrigerator, did you understand that?

OpenStudy (anonymous):

Eh, Kinda? :p

OpenStudy (anonymous):

@wmj259 can you help go through the step in solving it?

OpenStudy (wmj259):

Most certainly, but first? is that being divided by -3x^2y or -3x*2y?

OpenStudy (anonymous):

-3x^2y !

OpenStudy (wmj259):

Now that most certainly makes a HUGE difference.

OpenStudy (wmj259):

Be careful with exponents :).

OpenStudy (anonymous):

When I divided them all I got (not sure if its correct)